# Scalable Bayesian Learning with posteriors

### v1

@@ Appendix C SGMCMC for optimization
Let us consider gradient descent (with momenta) [72] (parameterization from PyTorch) for minimising $U(\theta)$:
@@
We see an equivalence between SGD and SGHMC ($T=0$) [77] with
@@
We note that the transition from SGMCMC to optimization and back is not ubiquitous amongst methods. Indeed, the Euler-Maruyama SGNHT sampling algorithm B.3, does not result in a functional optimization algorithm at the zero temperature limit as the value thermostat can only increase and results in divergant dynamics. Also, the popular Adam [39] optimizer does not fit the general sampling SDE (9) due to its use of elementwise operations, which do not have a convenient matrix-vector-product form. posteriors’ framework makes these mathematical considerations transparent and an easily accessible avenue for research.
@@ Appendix D Experiment details: Cold posterior effect
We now give additional details for the experiment into generalization capabilities of the Bayesian methods in posteriors and the cold posterior effect, from 5.1. Fully reproducible code can be found at posteriors/examples/imdb.
We here display more details on the results displayed in Figure 4. Figure 8 is exactly the same as Figure 4, but with the test loss for the MLE, approach added (no prior), which severely overfits. Figure 9 reports all accuracies which agree with the takeaways discussed in 5.1.
Figure 8: Cold posterior test losses (lower is better) for a range of posteriors algorithms and temperatures for the CNN-LSTM model [77] on the IMDB dataset [49] from 5.1 (same as Figure 4 with MLE added). Laplace EF is indistinguishable from Laplace GGN.
Figure 9: Cold posterior test accuracies (higher is better) for a range of posteriors algorithms and temperatures for the CNN-LSTM model [77] on the IMDB dataset [49] from 5.1.
We use exactly the same model CNN-LSTM from [77, 35], which consists of an embedding layer to convert input sequences into dense vectors, a 1D convolutional layer to capture local features, ReLU activation and max-pooling layers, an LSTM layer to capture long-term dependencies, and a fully connected layer to produce the final classification logits. The model has 2.7 million trainable parameters. Following [35] we use a diagonal Gaussian prior with all variances set to 1/40 (aside from MLE which has no prior or equivalently infinite prior variances).
We train on the IMDB [49] dataset for binary classification of positive/negative sentiment. We follow the default 50-50 split for the dataset with 25 thousand samples for training and 25 thousand samples for testing. All methods used batch size 32.
@@
  • MLE: We train using the AdamW optimizer [47] with all hyperparameters set to the detaults from TorchOpt [66] (learning rate $10^{-3}$).
@@
  • Variational Inference: As above, we train using the AdamW optimizer [47] with all default parameters. We apply the stick-the-landing variance reduction [70] and parameterization [67] tricks. We use a single sample Monte Carlo estimate at each step. We train the variational variances in log space to avoid negative variances and initialise all log standard deviations to $-3$. For test set evaluation, we sample 50 samples in parameter space and $10^{4}$ samples in logit space for the linearized approach.
@@ Appendix E Experiment details: Continual learning with LoRA
In this section, we detail details for the continual learning experiment from 5.2. Fully reproducible code can be found at posteriors/examples/continual_lora.
Our dataset [65] for the continual learning experiment, a collection of long books, is divided into $N$ episodes of train and test data. In the results reported in Figures 6 and 6, we use 1 book per episode, holding out the last $15\%$. We select the first twenty books from [65] that satisfy the following criterion:
@@
We use LoRA [31] on Llama 2 [74] to reduce the dimension of the inference task to 200k trainable parameters. We test two baseline methods. We fine-tune the model on each episode’s new data, a standard approach known to result in catastrophic forgetting, and we also implement a static offline baseline that sees all data every episode. This represents the LoRA network’s total capacity but is computationally infeasible for online learning.
The prior for the next episode is the posterior from the previous episode, becoming a quadratic penalty in the loss function during gradient descent. Whereas the original work on catastrophic forgetting [40] suggested using multiple penalties (incorporating data from all previous episodes into the prior), we use a single penalty following [32]. We use only the last episode’s posterior, reminiscent of exact Bayesian updates.
@@
    We use rank $r=8$ and $\alpha=32$ for LoRA, the standard settings [31], and fine-tune three weight matrices in the last layer (key, query, and output projections), following the literature [81].
@@ Appendix F Experiment details: Bayesian Llama 3
Here we describe further details for our final experiment fine-tuning the 8B Llama 3 model [1]. Fully reproducible code can be found at posteriors/examples/bayes_llama3.
As previously mentioned, our experiments involve using 10 sets of ensemble models for the last attention layer of the Llama3 model, resulting in 218 million trainable parameters. We train via SGHMC sampling for 20 epochs over the TQA dataset [37]. For evaluation, we generated a random set of 100 scientific factual statements from ChatGPT, which we also validated by hand. Additional hyperparameter and training details are as follows:

### v2

@@ Appendix C SGMCMC for optimization
Let us consider gradient descent (with momenta) [Sutskever et al., 2013], parameterization from PyTorch, for minimising $U(\theta)$:
@@
We see an equivalence between SGD and SGHMC ($T=0$) [Wenzel et al., 2020] with
@@
We note that the transition from SGMCMC to optimization and back is not ubiquitous amongst methods. Indeed, the Euler-Maruyama SGNHT sampling algorithm B.3, does not result in a functional optimization algorithm at the zero temperature limit as the value thermostat can only increase and results in divergant dynamics. Also, the popular Adam [Kingma and Ba, 2014] optimizer does not fit the general sampling SDE equation 9 due to its use of elementwise operations, which do not have a convenient matrix-vector-product form. posteriors’ framework makes these mathematical considerations transparent and an easily accessible avenue for research.
@@ Appendix D Experiment details: Cold posterior effect
We now give additional details for the experiment into generalization capabilities of the Bayesian methods in posteriors and the cold posterior effect, from Section 5.1. Fully reproducible code can be found at posteriors/examples/imdb.
We here display more details on the results displayed in Figure 4. Figure 8 is exactly the same as Figure 4, but with the test loss for the MLE, approach added (no prior), which severely overfits. Figure 9 reports all accuracies which agree with the takeaways discussed in Section 5.1.
Figure 8: Cold posterior test losses (lower is better) for a range of posteriors algorithms and temperatures for the CNN-LSTM model [Wenzel et al., 2020] on the IMDB dataset [Maas et al., 2011] from Section 5.1 (same as Figure 4 with MLE added). Laplace EF is indistinguishable from Laplace GGN.
Figure 9: Cold posterior test accuracies (higher is better) for a range of posteriors algorithms and temperatures for the CNN-LSTM model [Wenzel et al., 2020] on the IMDB dataset [Maas et al., 2011] from Section 5.1.
We use exactly the same model CNN-LSTM from Wenzel et al. [2020], Izmailov et al. [2021], which consists of an embedding layer to convert input sequences into dense vectors, a 1D convolutional layer to capture local features, ReLU activation and max-pooling layers, an LSTM layer to capture long-term dependencies, and a fully connected layer to produce the final classification logits. The model has 2.7 million trainable parameters. Following Izmailov et al. [2021] we use a diagonal Gaussian prior with all variances set to 1/40 (aside from MLE which has no prior or equivalently infinite prior variances).
We train on the IMDB Maas et al. [2011] dataset for binary classification of positive/negative sentiment. We follow the default 50-50 split for the dataset with 25 thousand samples for training and 25 thousand samples for testing. All methods used batch size 32.
@@
  • MLE: We train using the AdamW optimizer [Loshchilov and Hutter, 2018] with all hyperparameters set to the detaults from TorchOpt [Ren et al., 2023] (learning rate $10^{-3}$).
@@
  • Variational Inference: As above, we train using the AdamW optimizer [Loshchilov and Hutter, 2018] with all default parameters. We apply the stick-the-landing variance reduction [Roeder et al., 2017] and parameterization [Rezende et al., 2014] tricks. We use a single sample Monte Carlo estimate at each step. We train the variational variances in log space to avoid negative variances and initialise all log standard deviations to $-3$. For test set evaluation, we sample 50 samples in parameter space and $10^{4}$ samples in logit space for the linearized approach.
@@ Appendix E Experiment details: Continual learning with LoRA
In this section, we detail details for the continual learning experiment from Section 5.2. Fully reproducible code can be found at posteriors/examples/continual_lora.
Our dataset Rae et al. [2019] for the continual learning experiment, a collection of long books, is divided into $N$ episodes of train and test data. In the results reported in Figures 6 and 6, we use 1 book per episode, holding out the last $15\%$. We select the first twenty books from Rae et al. [2019] that satisfy the following criterion:
@@
We use LoRA [Hu et al., 2021] on Llama 2 [Touvron et al., 2023] to reduce the dimension of the inference task to 200k trainable parameters. We test two baseline methods. We fine-tune the model on each episode’s new data, a standard approach known to result in catastrophic forgetting, and we also implement a static offline baseline that sees all data every episode. This represents the LoRA network’s total capacity but is computationally infeasible for online learning.
The prior for the next episode is the posterior from the previous episode, becoming a quadratic penalty in the loss function during gradient descent. Whereas the original work on catastrophic forgetting [Kirkpatrick et al., 2017] suggested using multiple penalties (incorporating data from all previous episodes into the prior), we use a single penalty following Huszár [2018]. We use only the last episode’s posterior, reminiscent of exact Bayesian updates.
@@
    We use rank $r=8$ and $\alpha=32$ for LoRA, the standard settings [Hu et al., 2021], and fine-tune three weight matrices in the last layer (key, query, and output projections), following the literature [Yang et al., 2024].
@@ Appendix F Experiment details: Bayesian Llama 3
Here we describe further details for our final experiment fine-tuning the 8B Llama 3 model [AI@Meta, 2024]. Fully reproducible code can be found at posteriors/examples/bayes_llama3.
As previously mentioned, our experiments involve using 10 sets of ensemble models for the last attention layer of the Llama3 model, resulting in 218 million trainable parameters. We train via SGHMC sampling for 20 epochs over the TQA dataset [Kembhavi et al., 2017]. For evaluation, we generated a random set of 100 scientific factual statements from ChatGPT, which we also validated by hand. Additional hyperparameter and training details are as follows:
@@
Appendix G Bayesian logistic regression: Comparison and composition with Pyro
In this section, we use the probabilistic programming language (PPL) Pyro [Bingham et al., 2019] to construct a more classical Bayesian logistic regression model on the Pima Indians dataset [Smith et al., 1988] which is a standard benchmark in computational statistics, see e.g. Titsias [2024]. The model is constructed in Pyro and seamlessly converted into a log_posterior function for scalable inference with posteriors. These simulations demonstrate appropriate convergence of the posteriors in a smaller-scale setting, compare efficiency to other frameworks and demonstrate seamless composition with a PPL in Pyro.
We compare posteriors’ SGHMC and VI implementations with Pyro’s automated NUTS [Hoffman et al., 2014] implementation (which represents a gold standard) and a JAX [Bradbury et al., 2018] implementation with BlackJAX [Cabezas et al., 2024]. All serial MCMC methods were run for 5000 samples with similar warmup and thinning configurations (noting that NUTS has an adaptive thinning procedure). VI was ran for 8000 iterations which was sufficient for convergence. For parallel SGHMC we run 5000 samples independently for 2000 iterations.
We observe, in Figure 11, that all methods (including mini-batched versions, Figure 12) are broadly in agreement and have roughly converged to the same posterior distribution. In Table LABEL:tab:PI, we compare the fidelity of the posterior approximations by visualising the marginal distributions and comparing quantitatively via a kernelized Stein discrepancy [Gorham and Mackey, 2015]. We also display the runtimes of the approaches although in this case the model is small (8 dimensions and 768 data points) and therefore the Python overheads are relatively significant over model calls (as demonstrated by only a marginal speedup from mini-batching) and this is more so for PyTorch over JAX (although this rapidly deteriorates in the case of larger, more expensive model calls, in our experience). Nevertheless, we see that posteriors is competitive with Pyro [Bingham et al., 2019]. Importantly, these simulations demonstrate that posteriors can easily be composed one of the most popular probabilistic programming languages in Pyro [Bingham et al., 2019] to perform efficient and scalable mini-batch inference in traditional Bayesian models.
Figure 11: Full batch marginal distributions for Bayesian logistic regression on the Pima Indians dataset. All methods generated 5000 samples and provided approximations that are broadly in agreement.
Figure 12: Mini batch marginal distributions for Bayesian logistic regression. Otherwise, same setup as Figure 11 with full batch Pyro retained as a baseline. Again all methods provide relatively consistent posterior approximations.
Table 3: Metrics for Bayesian Logistic Regression. Kernelized Stein discrepancy (KSD) Gorham and Mackey [2015] measures the distance between the samples provided by the algorithm and the true posterior via a kernel function (in this case a standard Gaussian). All results are averaged over 10 random seeds with one standard deviation displayed. The displayed parallel SGHMC time represents the time for a single chain that could be obtained with sufficient parallel resources.
KSD $\downarrow$ Time (s) $\downarrow$
Pyro (NUTS Hoffman et al. [2014]) 2.680 ± 0.001 15.38 ± 0.74
BlackJAX (SGHMC, Full Batch) 2.689 ± 0.013 5.29 ± 0.74
posteriors (VI, Full Batch) 2.623 ± 0.001 7.04 ± 0.23
posteriors (VI, Batch Size=32) 2.516 ± 0.003 6.52 ± 0.22
posteriors (SGHMC, Full Batch) 2.693 ± 0.024 14.37 ± 0.31
posteriors (SGHMC, Batch Size=32) 2.686 ± 0.012 13.73 ± 0.38
posteriors (Parallel SGHMC, Batch Size=32) 2.690 ± 0.001 0.79 ± 0.04